### v1

@@ 2.1 Parameterized Complexity
Let $\Sigma$ be a finite alphabet. A classical problem $\mathcal{P}$ is a subset of $\Sigma^{*}$. Usually $\mathcal{P}$ consists of proper encodings of some combinatorial objects, e.g., graphs. A parameterized problem is a subset $\mathcal{P}\in\Sigma^{*}\times\mathbb{N}$. For an input $(x,k)\in\Sigma^{*}\times\mathbb{N}$, the second part $k$ is called the parameter.
@@
On instances where the parameter is relatively small, FPT and XP algorithms are efficient. In the case of diverse problems, we are usually interested in finding a small set of diverse solutions, so the number of requested solutions is a natural candidate for parameterization.
@@ 2.2 Cliquewidth
  • $t=intro(v,a)$, where $a\in[\omega]$ and $t$ is a leaf of $\mathcal{D}$. This node corresponds to construing one-vertex colored graph with a vertex $v$ of color $a$.
  • $t=t_{1}\oplus t_{2}$, where $t_{1},t_{2}$ are children of $t$. This node corresponds to taking disjoint union of colored graphs constructed by the subtrees of $t_{1},t_{2}$.
@@
  • $t=addEdges(t^{\prime},a,b)$, where $t^{\prime}$ is the only child of $t$, $a,b\in[\omega]$. This node corresponds to adding an edge between each pair of vertices $u,v$ of colors $a,b$ respectively in a graph constructed by subtree of $t^{\prime}$ if such edge does not exist.
For a node $t\in V(D)$ we denote by $\mathcal{D}_{t}$ the subtree of $\mathcal{D}$ rooted at $t$, by $G_{t}$ the colored graph constructed by $\mathcal{D}_{t}$, and by $\delta(t)$ the number of children of $t$.
Let $G$ be a simple, undirected graph. $\mathcal{D}$ is a cliquewidth decomposition of $G$ if $G$ is isomorphic to $G_{t}$, where $t$ is the root of $\mathcal{D}$.
Observe that in this definition for each $v\in V(G)$ there is exactly one node $t=intro(v,a)$, where $a\in[\omega]$. We also write $t=Leaf_{\mathcal{D}}(v)$. The cliquewidth of $G$ is minimal $\omega\in\mathbb{N}$ such that $G$ admits a decomposition $\mathcal{D}$ of width $\omega$.
@@ 2.3 $DivSum$ Is a Venn $f$-diversity Measure
where $\gamma(X,Y,v)=1$ if $v\in X\setminus Y\cup Y\setminus X$, 00 otherwise. Intuitively, $\gamma(X,Y,v)$ is $1$ if and only if the element $v$ contributes to the Hamming distance between $X$ and $Y$. Observe that
$\sum_{1\leq i<j\leq r}\gamma(S_{i},S_{j},v)$ $=\sum_{1\leq i<j\leq r}[v\in S_{i}\setminus S_{j}]+\sum_{1\leq i<j\leq r}[v\in S_{j}\setminus S_{i}]=$
$=\sum_{1\leq i<j\leq r}[v\in S_{i}\setminus S_{j}]+\sum_{1\leq j<i\leq r}[v\in S_{i}\setminus S_{j}]=$
@@ 2.4 Graph Logic
Graph problems can be described using logical formulas. In the first-order logic formula, only quantification over single variables is allowed, while in the second-order logic, we allow quantification over predicates. $MSO$ is the fragment of second-order logic where the second-order quantification is limited to unary predicates, which can be also viewed as subsets of the universe.
A graph can be modeled as a logical structure in several ways. The universe may be a set of vertices and the signature may consist of one binary relation, the adjacency. This model is called $MSO_{1}$.
Another way is to use vertices and edges as the universe, with the binary incidence relation. This model is denoted as $MSO_{2}$.
It can be shown that allowing quantification over sets of edges in $MSO_{2}$ makes this model more powerful than $MSO_{1}$. For example, the existence of Hamiltonian Path can be expressed in $MSO_{2}$, but not in $MSO_{1}$. However, we will focus on the first model.
Definition 2.4.
@@
such that $G\in\mathcal{P}$ if and only if $G\models\Phi$.
Similarly, an $MSO_{1}$ formula for a vertex problem $\mathcal{P}$ is such $\Phi$ (with one free variable) that $(G,S)\in\mathcal{P}$ if and only if $G\models\Phi(S)$.
3 Example: The Diverse Vertex Cover Problem
In the Vertex Cover problem given a graph $G$ and integer $k$, one asks whether there exists $S\subseteq V(G)$ of size at most $k$ such that $S$ touches all the edges of $G$. Formally, a k-Vertex Cover is a set of pairs $(G,S)$ where $S\subseteq V(G)$ such that $|S|\leq k$ and for each $uv\in E(G)$ $u\in S$ or $v\in S$.
In the diverse version of this problem, given a graph $G$ and integers $k,r,d$, one asks whether there exist $r$ vertex covers in $G$, each of size at most $k$, such that their diversity is at least $d$.
We will start with showing a dynamic programming that, given a cliquewidth decomposition $\mathcal{D}$ of a graph $G$ of width $\omega$, decides whether there exists a k-Vertex Cover of $G$ in time $f(\omega,k)\cdot|V(\mathcal{D})|$.
3.1 Dynamic Programming for Vertex Cover
Let $\mathcal{D}$ be a cliquewidth decomposition of $G$ of width $\omega$. For $t\in V(\mathcal{D})$ let $\mathcal{F}_{t}$ be a set of tuples $(c_{1},\ldots,c_{\omega})$ such that in $G_{t}$ there exist a vertex cover $S$ of size at most $k$ that uses $c_{i}$ vertices of color $i$ for each $i\in[\omega]$. Note that $|\mathcal{F}_{t}|\leq(k+1)^{\omega}$.
  1. 1.
    $t=intro(v,a)$
    $\mathcal{F}_{t}=\{C_{\{v\}},C_{\emptyset}\}$
    and $\mathcal{F}_{t}$ can be computed in time $2\omega$.
  2. 2.
    $t=t_{1}\oplus t_{2}$
    Observe that a set $S$ is a vertex cover in $G_{t}$ if and only if it is a union of vertex covers $S_{1}$ in $G_{t_{1}}$ and $S_{2}$ in $G_{t_{2}}$.
    Denote $(x_{1},\ldots,x_{\omega})+(y_{1},\ldots,y_{\omega})=(x_{1}+y_{1},\ldots,x_{\omega}+y_{\omega})$ and $|(x_{1},\ldots,x_{\omega})|=x_{1}+\ldots+x_{\omega}$. Then
    $\mathcal{F}_{t}=\{C_{1}+C_{2}|C_{1}\in\mathcal{F}_{t_{1}},C_{2}\in\mathcal{F}_{t_{2}},|C_{1}+C_{2}|\leq k\}$
    and $\mathcal{F}_{t}$ can be computed in time $(k+1)^{2\omega}\cdot\omega$.
  3. 3.
    $t=recolor(t^{\prime},a\rightarrow b)$
    Then
    $\mathcal{F}_{t}=\{\rho_{a\rightarrow b}(C^{\prime})|C^{\prime}\in\mathcal{F}_{t^{\prime}}\}$
    and $\mathcal{F}_{t}$ can be computed in time $(k+1)^{\omega}\cdot 2$.
  4. 4.
    $t=addEdges(t^{\prime},a,b)$
    Observe that a set $S$ is a vertex cover in $G_{t}$ if and only if it is a vertex cover in $G_{t^{\prime}}$ and contains all vertices of color $a$ or all vertices of color $b$. Let $n_{i}$ be the number of vertices of color $i$ in $G_{t}$. We can compute it by induction for each $i\in[\omega]$ for each $t\in V(\mathcal{D})$ in time $|V(\mathcal{D})|$. Then
    and $\mathcal{F}_{t}$ can be computed in time $(k+1)^{\omega}\cdot 2$.
@@ 3.1 Dynamic Programming for Vertex Cover
$\mathcal{O}(|V(\mathcal{D})|\cdot(k+1)^{2\omega}\cdot\omega).$
@@
For each node $t\in V(\mathcal{D})$ set $\mathcal{F}_{t}$ can be computed in time
$max(2\omega,(k+1)^{2\omega}\cdot\omega,(k+1)^{\omega}\cdot 2).$
3.2 From Vertex Cover to Diverse Vertex Cover
We will start with proving the lemma that will let us compute the diversity of the tuple of partial solutions $(S_{1},\ldots,S_{r})$ for $t\in V(\mathcal{D})$ based on the diversity values of tuples of partial solutions for the children of $t$.
@@ 3.2 From Vertex Cover to Diverse Vertex Cover
$Div(S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r})=Div(S_{1},\ldots,S_{r})+Div(P_{1},\ldots,P_{r})-|V|\cdot f(0^{r}).$
Proof.
@@
where $m(S_{1},\ldots,S_{r},v)$ is the membership vector of $v$ in $S_{1},\ldots,S_{r}$. If $v\not\in S$, $m(S_{1},\ldots S_{r},v)=0^{r}$, and analogously if $v\not\in P$, so we get
$Div(S_{1},\ldots,S_{r})=\sum_{v\in V}f(m(S_{1},\ldots,S_{r},v))=\sum_{v\in S}f(m(S_{1},\ldots,S_{r},v))+|P|\cdot f(0^{r})+|V\setminus(S\cup P)|\cdot f(0^{r}),$
$Div(P_{1},\ldots,P_{r})=\sum_{v\in V}f(m(P_{1},\ldots,P_{r},v))=\sum_{v\in P}f(m(P_{1},\ldots,P_{r},v))+|S|\cdot f(0^{r})+|V\setminus(S\cup P)|\cdot f(0^{r}).$

### v2

@@ 2.1 Parameterized Complexity
Let $\Sigma$ be a finite alphabet. A classical problem $\mathcal{P}$ is a subset of $\Sigma^{*}$. Usually, $\mathcal{P}$ consists of proper encodings of some combinatorial objects, e.g., graphs (then we call $\mathcal{P}$ a graph problem). A parameterized problem is a subset $\mathcal{P}\in\Sigma^{*}\times\mathbb{N}$. For an input $(x,k)\in\Sigma^{*}\times\mathbb{N}$, the second part $k$ is called the parameter.
@@
In instances where the parameter is relatively small, FPT and XP algorithms are efficient. In the case of diverse problems, we are usually interested in finding a small set of diverse solutions, so the number of requested solutions is a natural candidate for parameterization.
@@ 2.2 Cliquewidth
  • $t=intro(v,a)$, where $a\in[\omega]$ and $t$ is a leaf of $\mathcal{D}$. This node corresponds to construing a one-vertex colored graph with a vertex $v$ of color $a$.
  • $t=t_{1}\oplus t_{2}$, where $t_{1},t_{2}$ are children of $t$. This node corresponds to taking the disjoint union of colored graphs constructed by the subtrees of $t_{1},t_{2}$.
@@
  • $t=addEdges(t^{\prime},a,b)$, where $t^{\prime}$ is the only child of $t$, $a,b\in[\omega]$. This node corresponds to adding an edge between each pair of vertices $u,v$ of colors $a,b$ respectively in a graph constructed by a subtree of $t^{\prime}$ if such an edge does not exist.
For a node $t\in V(\mathcal{D})$ we denote by $\mathcal{D}_{t}$ the subtree of $\mathcal{D}$ rooted at $t$, by $G_{t}$ the colored graph constructed by $\mathcal{D}_{t}$, and by $\delta(t)$ the number of children of $t$.
Let $G$ be a simple graph. $\mathcal{D}$ is a cliquewidth decomposition of $G$ if $G$ is isomorphic to $G_{t}$ for $t$ root of $\mathcal{D}$.
Observe that in this definition, for each $v\in V(G)$, there is exactly one node $t=intro(v,a)$, where $a\in[\omega]$. We also write $t=Leaf_{\mathcal{D}}(v)$. The cliquewidth of $G$ is minimal $\omega\in\mathbb{N}$ such that $G$ admits a decomposition $\mathcal{D}$ of width $\omega$.
@@ 2.3 $DivSum$ is a Venn $f$-diversity Measure
where $\gamma(X,Y,v)=1$ if $v\in X\setminus Y\cup Y\setminus X$, 0 otherwise. Intuitively, $\gamma(X,Y,v)$ is $1$ if and only if the element $v$ contributes to the Hamming distance between $X$ and $Y$. Observe that
$\sum_{1\leq i<j\leq r}\gamma(S_{i},S_{j},v)$ $=\sum_{1\leq i<j\leq r}[v\in S_{i}\setminus S_{j}]+\sum_{1\leq i<j\leq r}[v\in S_{j}\setminus S_{i}]$
$=\sum_{1\leq i<j\leq r}[v\in S_{i}\setminus S_{j}]+\sum_{1\leq j<i\leq r}[v\in S_{i}\setminus S_{j}]$
@@ 2.4 Graph Logic
Graph problems can be described using logical formulas. In the first-order logic formula, only quantification over single variables is allowed, while in the second-order logic, we allow quantification over predicates. $MSO$ is the fragment of second-order logic where the second-order quantification is limited to unary predicates, which can also be viewed as subsets of the universe. A graph can be modeled as a logical structure in several ways. The universe may be a set of vertices, and the signature may consist of one binary relation, the adjacency. This model is called $MSO_{1}$. Another way is to use vertices and edges as the universe, with the binary incidence relation. This model is denoted as $MSO_{2}$. It can be shown that allowing quantification over sets of edges in $MSO_{2}$ makes this model more powerful than $MSO_{1}$. For example, the existence of a Hamiltonian Path can be expressed in $MSO_{2}$, but not in $MSO_{1}$. See, for example, [CE12] for further details. In this paper, we focus on the first model.
Definition 2.4 ($MSO_{1}$ graph logic).
@@
such that $G\in\mathcal{P}$ if and only if $G\models\Phi$. Similarly, an $MSO_{1}$ formula for a vertex problem $\mathcal{P}$ is $\Phi$ with one free variable such that $(G,S)\in\mathcal{P}$ if and only if $G\models\Phi(S)$.
3 Monotone DP for Vertex Problems
A very natural property of dynamic programming algorithms for vertex problems on cliquewidth decompositions is that each partial solution for a node $t$ is a disjoint union of partial solutions for its children. We call algorithms with this property monotone.
In this section, we give the formalism of dynamic programming and the monotone property. We prove our main result, that having a monotone dynamic program for each of the vertex problems $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ and a cliquewidth decomposition $\mathcal{D}$ of $G$, we can solve $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on $G$.
3.1 Definitions
Recall that $\delta(t)$ denotes the number of children of a node $t\in V(\mathcal{D})$.
Definition 3.1 (Dynamic programming core).
A dynamic programming core is an algorithm $\mathfrak{C}$ that takes cliquewidth decomposition $\mathcal{D}$ of a graph $G$ and produces
  • A finite set $Accept_{\mathfrak{C},\mathcal{D}}\subseteq\{0,1\}^{*}$,
Let $\tau(\mathfrak{C},\mathcal{D})$ be the time necessary to construct these data and $Size_{\mathfrak{C},\mathcal{D}}\coloneqq max_{t\in V(\mathcal{D})}|Process_{\mathfrak{C,\mathcal{D}}}(t)|$.
Intuitively, $Process_{\mathfrak{C},\mathcal{D}}$ corresponds to the transitions of the dynamic programming. If a partial solution represented by $w$ at node $t$ can be obtained from partial solutions represented by $w_{1}\ldots w_{\delta(t)}$ at its children $t_{1}\ldots t_{\delta(t)}$ then $ww_{1}\ldots w_{\delta(t)}\in Process_{\mathfrak{C},\mathcal{D}}(t)$.
Definition 3.2 (Witness).
  • for each $t\in V(\mathcal{D})$ with children $t_{1},\ldots,t_{\delta(t)}$, $\alpha(t)\alpha(t_{1})\ldots\alpha(t_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)$,
  • $\alpha(q)=w$.
Observe that if $\mathcal{P}$ is a vertex problem, given a graph $G$, we cannot hope to list all $S$ such that $(G,S)\in\mathcal{P}$ in FPT time, as the number of such sets may be exponential in $|G|$. Therefore, we focus on deciding whether for a given graph $G$ there exists any such $S$. We call this decision problem a corresponding graph problem and write $G\in\mathcal{P}$ if the answer is Yes.
Definition 3.3.
We say that a dynamic programming core $\mathfrak{C}$ solves a graph problem $\mathcal{P}$ if for each graph $G$ and for each cliquewidth decomposition $\mathcal{D}$ of $G$, $G\in\mathcal{P}$ if and only if a $(\mathfrak{C},\mathcal{D})$-witness exists.
@@ 3.1 Definitions
  • If $t$ has children $t_{1},\ldots t_{\delta(t)}$, $\Pi(t)=\{w:\exists_{(w,w_{1},\ldots w_{\delta(t)})\in Process_{\mathfrak{C},\mathcal{D}}(t)}\ s.t.\ \forall_{i\in[\delta(t)]}\ w_{i}\in\Pi(t_{i})\}$.
$G\in\mathcal{P}$ if and only if $\Pi(q)\cap Accept_{\mathfrak{C},\mathcal{D}}\not$$\emptyset$, where $q$ is the root of $\mathcal{D}$. ∎
Note that from the above algorithm we can also easily recover some $(\mathfrak{C},\mathcal{D},w)$-witness for each $w\in\Pi(q)$.
Definition 3.5 (Monotonicity).
Let $\mathcal{P}$ be a vertex problem. We call a dynamic programming core $\mathfrak{C}$ monotone if there exists a function $\rho:\{0,1\}^{*}\rightarrow\{0,1\}$ such that for each graph $G$, for each cliquewidth decomposition $\mathcal{D}$ of $G$, for each $S\subseteq V(G)$, $(G,S)\in\mathcal{P}$ if and only if there exists a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ such that $S=\{v\in V(G):\rho(\alpha(Leaf_{\mathcal{D}}(v)))=1\}$}. We call such $\rho$ a $\mathfrak{C}$-vertex-membership function and say a pair $(\mathfrak{C},\rho)$ solves $\mathcal{P}$.
Intuitively, a vertex membership function $\rho$ together with a cliquewidth decomposition $\mathcal{D}$ of a graph $G$ and a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ provide an encoding of a subset of vertices of $G$ using only the values of DP entries in leaves of $\mathcal{D}$. Let $S_{\rho}(\mathcal{D},\alpha)=\{v\in V(G):\rho(\alpha(Leaf_{\mathcal{D}}(v)))=1\}$. Note that in this definition, we determine whether a vertex is in the subset or not based only on the value of the witness on its introduce node.
3.2 Proof of Theorem˜1.2
The following theorem is a formal reformulation of Theorem˜1.2 stated in the introduction. Indeed, if $t_{\mathcal{D}}=\mathcal{O}(Size_{\mathfrak{C}_{i},\mathcal{D}})$ for $\mathcal{P}_{i}$ being the one with the slowest maximal running time on a single node of the decomposition out of $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$, and assuming we are given $Accept_{\mathfrak{C}_{i},\mathcal{D}}$ and $Process_{\mathfrak{C}_{i},\mathcal{D}}(t)$ for each $t\in V(\mathcal{D})$ precomputed, the time needed to solve $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ is
$\mathcal{O}\left(|V(\mathcal{D})|\cdot\prod_{i=1}^{r}Size_{\mathfrak{C}_{i},\mathcal{D}}\right)\leq\mathcal{O}(|V(\mathcal{D})|t_{\mathcal{D}}^{r}).$
Theorem 3.6.
Before proving it, we will also state the following theorem, which is a formal reformulation of Theorem˜1.3 from the introduction. The proof of Theorem˜3.6 can be easily modified to work for Theorem˜3.7.
Theorem 3.7.
Keep in mind that the following proof can be modified to work also for $DivMin$ as stated in Theorem˜3.7. Instead of one diversity value $\ell$ for a tuple of partial solutions, we need to store a vector $L\in\{0,\ldots,d\}^{\binom{r}{2}}$ of pairwise Hamming distances between the solutions. Since for $S_{1},S_{2}\subseteq S,\ P_{1},P_{2}\subseteq P$, where $S$ and $P$ are disjoint, $HamDist(S_{1}\cup P_{1},S_{2}\cup P_{2})=HamDist(S_{1},S_{2})+HamDist(P_{1},P_{2})$, we can sum these vectors point-wise as we sum single diversity values. As the Hamming distance only grows, and we are interested in having the minimum distance of at least $d$, we can treat all the values greater than $d$ as equal to $d$. However, we cannot take advantage of keeping only the best diversity value for a tuple of entries because such vectors may be incomparable, so for one tuple of entries that represents multiple tuples of partial solutions, there may be up to $(d+1)^{\binom{r}{2}}$ different diversity vectors, which contributes to $d^{r^{2}}$ factor in the running time of Theorem˜3.7.
The correctness of the algorithm is based on the following lemma and the fact that for each node $t=t_{1}\oplus t_{2}$ of the decomposition, $V(G_{t_{1}})\cap V(G_{t_{2}})=\emptyset$.
@@
$Div(S_{1}\cup P_{1},\ldots,S_{r}\cup P_{r})=Div(S_{1},\ldots,S_{r})+Div(P_{1},\ldots,P_{r})-|V(G)|\cdot f(0^{r}).$
Note that in Lemma˜3.8, we need to know the number of vertices of the final graph ($|V(G)|$). However, $f(0^{r})$ corresponds to the influence of a vertex that is in none of the partial solutions, which could be set to 0 in many convenient measures. In such cases, we can even ignore this mild constraint.
Proof of Lemma˜3.8.
@@
where $m(S_{1},\ldots,S_{r},v)$ is the membership vector of $v$ in $S_{1},\ldots,S_{r}$. If $v\not\in S$, $m(S_{1},\ldots S_{r},v)=0^{r}$, and analogously if $v\not\in P$, $m(P_{1},\ldots P_{r},v)=0^{r}$ so we get
$Div(S_{1},\ldots,S_{r})=\sum_{v\in V}f(m(S_{1},\ldots,S_{r},v))=\sum_{v\in S}f(m(S_{1},\ldots,S_{r},v))+|P|f(0^{r})+|V\setminus(S\cup P)|f(0^{r}),$
$Div(P_{1},\ldots,P_{r})=\sum_{v\in V}f(m(P_{1},\ldots,P_{r},v))=\sum_{v\in P}f(m(P_{1},\ldots,P_{r},v))+|S|f(0^{r})+|V\setminus(S\cup P)|f(0^{r}).$